Andrew, Brenda and Cammi all work part-time at a
supermarket. Cammi earns $20 more than Andrew and
Brenda earns $30 less than twice Andrew’s wage. If
their total combined wage is $400, find how much each of these workers earns.

Answer by JulietG(1812)

A + B + C = 400
C = A + 20
B = 2A - 30
Take the known values of C & B from the latter equations and substitute into the first equation.
A + B + C = 400
A + (2A-30) + (A+20) = 400
4A - 10 = 400
Add 10 to each side
4A = 410
Divide each side by 4
A = 102.5
If Andrew earns 102.5, then Cammi earns 122.50, and Brenda earns 205-30 or 175.
102.5 + 122.5 + 175 = 400